# TiRE-GAN: Task-Incentivized Generative Learning Models for Radiomap Estimation with Radio Propagation Model

### v1

@@ IV-A Experiment Setup
IV-A1 Data description
In this work, we utilize the RadiomapSeer dataset [10] for evaluation, which comprises an extensive set of 700 maps from a variety of metropolitan areas, including Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. Each map delineates a quadrant of $256\times 256$ square meters, depicted with a granularity of $256\times 256$ pixels. The edifices are encoded by assigning a pixel intensity of one, whereas open spaces are represented by a zero value. Furthermore, the precise locations of transmitters are encoded with a pixel intensity of one superimposed on a null background, facilitating accurate localization.
IV-A2 Method Comparison
To evaluate the performance of our TiRE-GAN, we compare with both state-of-the-art (SOTA) model-based methods, including Kriging [17], and data-driven approaches, including Matrix completion (MC) [18], AutoEncoder (AE) [11], RadioUnet (RU) [10] and conventional cGAN [15]. For the interpolation-based method, we reconstruct the radiomap with observations in the same radiomap. For the prediction-based method, we split the dataset as 4000/800/800 (70%/15%/15%) for training/validation/testing sets. In our TiRE-GAN, we set $\alpha=1$ and $\beta=0.1$ in Eq. (9).
@@ IV-B Overall performance
We first present the overall performance of different radiomap estimations under different sampling ratios (sr). In this work, we evaluate the performance based on the normalized radiomaps via MSE denoted by
$\text{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (13)
and normalized MSE (NMSE) calculated by
$\text{NMSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}\frac{(y_{i}-\hat{y}_{i})^{2}}{y_{i}^{2}},$ (14)
where $K=256^{2}$. Note that, since the linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain. The results are presented in Table I and II. From the results, our proposed TiRE-GAN always achieves superior performance against all other SOTA approaches. Specifically, with the assistance of our proposed radio depth map, our method can achieve a very high reconstruction accuracy even with a very small sampling ratio and insufficient training samples, which validates the effectiveness of the proposed method. The visualization results of different methods in 3 radiomap examples are shown in Fig. 4, where the sample ratio is 1%. From the visualization results, the radioUnet, cGAN, and TiRE-GAN have better performances than Kringing, MC, and AE in capturing the overall radio propagation patterns. Compared to radioUnet and cGAN, our TiRE-GAN has more explicit boundaries for the shadowing effects, resulting from the compensation from the TIN feedback. The numerical results, together with the visualization results demonstrate the efficiency of our proposed TiRE-GAN for radiomap estimation.
TABLE I: Test NMSE Comparison (×$\times 10^{-2}$)
sr (%) 0.03 0.1 0.5 1 3 5 10 Average
Krigin 24.5696 18.9280 14.2451 12.9608 14.7171 11.3540 10.9647 15.3913
RU 6.5421 5.1111 2.6374 2.2206 1.8192 1.5191 1.2704 2.8744
AE 54.7656 53.8440 45.6640 43.3302 26.0041 20.7167 14.3275 36.9503
AE deep 50.6597 48.8197 21.4272 16.2537 6.8252 6.5503 5.8179 156.3537
cGAN 11.5754 5.4805 2.1345 1.8301 1.3009 1.1894 1.0295 3.5057
MC 83.0969 74.7810 69.0133 66.9996 62.8215 56.3645 55.4314 66.9297
TiRE-GAN 1.9866 1.2608 0.7282 0.6084 0.5228 0.5119 0.4615 0.8686
TABLE II: Test MSE Comparison (×$\times 10^{-4}$)
sr (%) 0.03 0.1 0.5 1 3 5 10 Average
Krigin 188.7554 141.3767 110.7954 101.5476 91.9940 89.0443 86.0208 115.5340
RU 24.6458 19.3893 10.0221 8.4693 7.0856 5.9640 4.9978 11.5105
AE 230.6564 226.6141 194.4702 186.5724 113.0439 87.7205 65.0134 157.7272
AE deep 209.2621 201.4013 91.9589 70.0012 26.8177 26.4321 23.4719 92.7636
cGAN 45.8797 22.0908 8.4572 7.2467 5.0700 4.6442 4.0541 13.9203
MC 342.1892 307.9448 284.1937 275.9010 258.6960 232.1065 228.2640 275.6136
TiRE-GAN 7.0604 4.6453 2.7678 2.3472 2.0377 2.0015 1.8137 3.2390
@@ IV-C Robustness Against Noise
To evaluate the robustness of TiRE-GAN, we conduct the experiment, where a noisy input pattern is substituted for the original sparse observation channel by introducing additive Gaussian noise to each sampled value in the radiomap. To showcase the effectiveness of the proposed radio depth map and TIN blocks, we compare our TiRE-GAN with Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), and conditional GAN (cGAN) under different signal-to-noise ratios (SNR) as Fig. 5. Here, we consider a sampling ratio of 1% for all the test cases. From the results, our TiRE-GAN always has better performance. Moreover, the MSE of TiRE-GAN increases slower than those of other methods, especially compared to the high performance methods (RadioUNet, cGAN), demonstrating the robustness of the TiRE-GAN against noise.
Figure 5: The MSE (×$\times 10^{-4}$) of Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and TiRE-GAN of different SNRs under a sample ratio=1%.
@@ IV-D Outage Evaluation of Generated radiomap
To further evaluate the performance of reconstructed radiomaps in downstream tasks, we compare the performance of RM-assisted outage detection. Here, we input the generated radiomaps from different methods into the pre-trained outage detection networks. The numerical results of the whole datasets with sr=1% and the exemplary visualization results are shown in Table III and Fig. 6, respectively. From the results, the traditional model-based methods and conventional auto-encoder perform poorly in the outage detection. Our TiRE-GAN always has superior performance, which demonstrates the effectiveness of TIN in radiomap compensation for downstream tasks.
TABLE III: Outage Map MSE comparison
Method Kriging MC AE
MSE 0.078134 0.075765 0.049377
Method RU cGAN TiRE-GAN
MSE 0.028439 0.030352 0.026682
Figure 6: Illustration of Outage Map Prediction under sr=1%. We compared the outage map predicted by the generated radiomap from Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), and our TiRE-GAN. The corresponding radiomap and the ground truth outage map(GT) are shown on the right for reference. The result shows that TiRE-GAN has the best capability to predict the outage map with a clear boundary and range.
@@ V Conclusion
In this work, we introduce a task-incentivized generative learning model, namely TiRE-GAN for radiomap estimation. More specifically, to characterize the overall radio propagation behavior, we introduce a radio depth map as additional physics-embedded features as the condition to guide the data-driven generative models. To compensate for the radiomap estimation for downstream tasks, we introduce a task-incentivized network to provide feedback to predict the detailed features in the radiomap. From our experimental results, the proposed TiRE-GAN always achieves superior performance, even with insufficient training samples, which addresses the barriers of conventional learning approaches. Our numerical and visualization results demonstrate the effectiveness of the proposed radio depth map and TiRE-GAN. Future research endeavors will delve into the optimization of the generator’s reward mechanism. Another promising direction is the multi-band radiomap estimation for different frequencies.

### v3

@@ IV-A Experiment Setup
IV-A1 Data
In this work, we utilize the RadiomapSeer dataset [8] for evaluation, which contain an extensive set of 700 maps from a variety of metropolitan areas, including Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. Each map delineates a quadrant of $256\times 256$ square meters, at a granularity of $256\times 256$ pixels. The edifices are encoded by assigning a pixel intensity of one, whereas open spaces are represented by a zero value. Furthermore, the precise locations of transmitters are encoded with a pixel intensity of one superimposed on a null background, facilitating accurate localization.
IV-A2 Comparison
For performance assessment, we compare TiRE-GAN with both state-of-the-art (SOTA) model-based methods, including Kriging [17], and data-driven approaches, including AutoEncoder (AE) [9], RadioUnet (RU) [8], conventional cGAN [15], RME-GAN [10], and Geo2SigMap [18]. For interpolation-based method, we reconstruct the radiomap with observations in the same radiomap. For prediction-based method, we split the dataset as 4000/800/800 (70%/15%/15%) for training/validation/testing sets. Since RME-GAN is designed for non-uniformly distributed samples, we remove the corresponding loss terms to ensure fair comparison. In our TiRE-GAN, we set $\alpha=1$ and $\beta=0.1$ in Eq. (8).
@@ IV-B Overall performance
We first present the overall performance of different radiomap estimations. We evaluate the performance based on the normalized radiomaps via MSE denoted by $\text{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(y_{i}-\hat{y}_{i})^{2}$, and normalized MSE (NMSE) calculated by $\text{NMSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\frac{(y_{i}-\hat{y}_{i})^{2}}{y_{i}^{2}}$, where $N_{p}=256^{2}$. Linear normalization makes it easy to re-scale the estimated radiomap back to the original data domain. From the results given in Table I, our proposed TiRE-GAN always achieves superior performance against all other approaches. Different from RME-GAN and purely data-driven models, our radio depth map can capture both pathloss and shadowing efficiently, providing efficient guidance of cGAN training. With the assistance of radio depth map, our method can achieve a very high reconstruction accuracy even for very low sampling ratio and few training samples. Fig. 3 presents the visualization results of different methods for three radiomap examples at a 1% sample ratio. Advanced learning methods, such as RadioUnet, cGAN, and our TiRE-GAN, outperform Kringing and AE in capturing overall radio propagation patterns. Notably, due to the model-based interpolation block and complex loss design, RME-GAN cannot efficiently function at very low sampling ratio below 0.1%. TiRE-GAN shows clearer boundaries for shadowing effects resulted from the TIN feedback. Both the visualization and numerical results validate the effectiveness of our TiRE-GAN in RME with insufficient training samples.
TABLE I: Test NMSE (×$\times 10^{-2}$)/MSE (×$\times 10^{-4}$) Comparison
sr (%) 0.03 0.1 0.5 1 3 5 10 Average ($0.5\sim 10$)
Kriging 24.56/188.75 18.92/141.37 14.24/110.79 12.96/101.54 14.71/91.99 11.35/89.04 10.96/86.02 12.84/95.88
AE 54.76/230.65 53.84/226.61 45.66/194.47 43.33/186.57 26.00/113.04 20.71/87.72 14.32/65.01 30.00/129.36
AE Deep 50.65/209.26 48.81/201.40 21.42/91.95 16.25/70.00 6.82/26.81 6.55/26.43 5.81/23.47 11.22/47.73
RU 6.54/24.64 5.11/19.38 2.63/10.02 2.22/8.46 1.81/7.08 1.51/5.96 1.27/4.99 1.89/7.30
cGAN 11.57/45.88 5.48/22.09 2.13/8.46 1.83/7.25 1.30/5.07 1.18/4.64 1.02/4.05 1.49/5.89
RME-GAN -/- -/- 0.77/7.42 0.68/5.70 0.64/6.15 0.53/5.08 0.47/4.51 0.60/5.77
Geo2SigMap 2.19/8.67 1.23/4.73 1.12/4.31 0.75/2.89 0.79/2.93 0.79/2.94 0.95/3.75 0.93/3.36
TiRE-GAN 1.98/7.06 1.26/4.65 0.72/2.77 0.61/2.35 0.52/2.04 0.51/2.00 0.46/1.81 0.55/2.19
@@ IV-C Robustness Against Noise
To assess the robustness, we include additive Gaussian noise to each sampled radiomap value, replacing the original sparse observation channel with a noisy input pattern. To highlight the effectiveness of the proposed radio depth map and TIN blocks, we compared different methods across various signal-to-noise ratios (SNR), as shown in Fig. 4, using a 1% sampling ratio for all test cases. TiRE-GAN consistently outperforms the other methods. Moreover, its MSE growth is slow, particularly compared to advanced learning methods like RadioUNet and cGAN, which further demonstrates robustness of TiRE-GAN against noise.
Figure 4: The MSE (×$\times 10^{-4}$) comparison with different SNRs.
@@ IV-D Outage Evaluation of Generated radiomap
To further evaluate the effect on downstream tasks, we compare the performance of RM-assisted outage detection. Here, we input the generated radiomaps from different methods into the pre-trained outage detection networks. The numerical results of the whole datasets with 1% sampling and the exemplary visualization results are shown in Table II and Fig. 5, respectively. RME-GAN and TiRE-GAN achieve lower MSE than other methods, resulted from their utilization of model information. However, in the visualization, TiRE-GAN displays smoother and more consistent patterns, which should benefit the spectrum management. These results demonstrate the effectiveness of TIN in radiomap compensation.
TABLE II: Outage Map MSE comparison
Method Kriging AE AEdeep RU
MSE 0.0781 0.0758 0.0494 0.0284
Method cGAN Geo2SigMap RME-GAN TiRE-GAN
MSE 0.0304 0.0561 0.0231 0.0267
Figure 5: Illustration of Outage Map Prediction under sr=1%. We compared the outage map predicted by the generated radiomap from different methods. The corresponding radiomap and the ground truth outage map(GT) are shown on the right for reference.
@@ V Conclusion
TiRE-GAN is a task-incentivized generative learning model for radiomap estimation. To capture the overall radio propagation behavior, it uses a radio depth map as a physics-embedded feature to guide the data-driven generative model. Additionally, a task-incentivized network provides feedback to enhance detailed radiomap features for downstream tasks. Experimental results show that TiRE-GAN consistently outperforms other methods, especially with limited training samples, a faced by realistic applications. Future work will focus on optimizing the generator’s reward mechanism and exploring multi-band radiomap estimation across different frequencies.